# Domains: Guiding LLM Reasoning for Explainable Answers in High-Stakes Domains

### v1

@@ Abstract
Large Language Models (LLMs) are widely applied to downstream domains. However, current LLMs for high-stakes domain tasks, such as financial investment and legal QA, typically generate brief answers without reasoning processes and explanations. This limits usersā confidence in making decisions based on their responses. While original CoT shows promise, it lacks self-correction mechanisms during reasoning. This work introduces Domains, which enhances LLMsā reasoning capabilities on domain tasks through supervised fine-tuning and tree search. We construct CoT-stock-2k and CoT-legal-2k datasets for fine-tuning models that activate domain-specific reasoning steps based on their judgment. Additionally, we propose Selective Tree Exploration to spontaneously explore solution spaces and sample optimal reasoning paths to improve performance. We also introduce PROOF-Score, a new metric for evaluating domain modelsā explainability, complementing traditional accuracy metrics with richer assessment dimensions. Extensive experiments on stock investment recommendation and legal reasoning QA tasks demonstrate Domainsās leading performance and explainability. Our code is available at https://anonymous.4open.science/r/Domaino1s-006F/.
@@ 7 Ethical Considerations
In this section, we discuss several important ethical considerations regarding the training, deployment, and use of Domains.
@@ 7.2 Potential Risks in Financial and Legal Applications
We implement safety checks in the Domain Safety (DS) metric of PROOF-Score to detect potentially harmful or high-risk advice.

### v2

@@ Abstract
Large Language Models (LLMs) are widely applied to downstream domains. However, current LLMs for high-stakes domain tasks, such as financial investment and legal QA, typically generate brief answers without reasoning processes and explanations. This limits usersā confidence in making decisions based on their responses. While original CoT shows promise, it lacks self-correction mechanisms during reasoning. This work introduces Domains, which enhances LLMsā reasoning capabilities on domain tasks through supervised fine-tuning and tree search. We construct CoT-stock-2k and CoT-legal-2k datasets for fine-tuning models that activate domain-specific reasoning steps based on their judgment. Additionally, we propose Selective Tree Exploration to spontaneously explore solution spaces and sample optimal reasoning paths to improve performance. We also introduce PROOF-Score, a new metric for evaluating domain modelsā explainability, complementing traditional accuracy metrics with richer assessment dimensions. Extensive experiments on stock investment recommendation and legal reasoning QA tasks demonstrate Domainsās leading performance and explainability. Our code is available at https://github.com/Hyalinesky/Domaino1s.
@@ 4.4 Ablation Study (RQ3)
5 Acknowledgment
This work is supported by the National Key R&D Program of China [2022YFF0902703].
@@ 8.2 Potential Risks in Financial and Legal Applications
We implement safety checks in the Domain Safety (DS) metric of PROOF-Score to detect potentially harmful or high-stake advice.
@@ Appendix H Reasoning Process Analysis
Appendix I Training Data Scaling
In this section, we explore the impact of increasing training data on Domains performance. We introduce more legal QA dataĀ Li etĀ al. (2022), and following the process in SectionĀ 3.2, we use GPT-4o to generate reasoning processes, expanding the CoT-legal dataset to 10k entries. Then, we train the Qwen-2.5-Instruction model with 2k, 5k, and 10k data separately, and evaluate the performance differences of Domains-legal trained with different training data volumes. The results are as follows:
| Model | CC | CAUD | MAUD | PP | IP | PJ | Scalar | TTP | TTD | Avg. |
| Domains-legal (2k) | 88.64 | 81.76 | 80.33 | 66.54 | 52.63 | 72.00 | 88.97 | 95.33 | 78.78 | 78.33 |
| Domains-legal (5k) | 88.73 | 81.74 | 80.65 | 66.95 | 51.88 | 74.00 | 89.32 | 93.93 | 77.57 | 78.30 |
| Domains-legal (10k) | 88.96 | 81.63 | 81.87 | 66.63 | 53.38 | 74.00 | 89.32 | 93.93 | 80.37 | 78.90 |
The results show that the accuracy improvement from increasing training data from 2k to 5k is not significant, and while increasing to 10k seems to increase the average accuracy (Avg.), the improvement on individual tasks is not substantial. This aligns with our approach of using a small number of training samples (2k) in this paper. The purpose of constructing a small number of training samples is not to introduce substantial domain-specific knowledge, but to stimulate the modelās reasoning abilities in specific domains. However, we also encourage applying our method to domain models trained with large-scale domain data, using our few-shot approach to activate their reasoning capabilities to achieve stronger performance and robustness.
Appendix J Human Evaluation Alignment
In SectionĀ 4.3, we introduce PROOF-Scores based on GPT-4o to automate the evaluation of modelsā reasoning capability. Considering that human evaluations may have inconsistent scales and aspects compared to LLMs, we encourage having a human expert in the loop in evaluation scenarios that do not require rapid response and automation, to avoid biases introduced by LLMs.
On the other hand, studiesĀ chiang2023can; zheng2023judging; desmond2024evalullm support that LLMs like GPT-4 have evaluation capabilities comparable to humans. To understand whether this also applies to high-stake domain tasks, and to determine if our PROOF-Scores (GPT-4o) align with human judgment, we invited 3 PhD in finance and 3 in law to score the explainability of model results in TableĀ 4 according to the scoring rules and intervals designed for PROOF-Scores. Given limited resources, we randomly sampled only 30% of questions from each dataset to create the questionnaires.
For expert scoring, we calculate the PROOF-Score the same way as (3). We calculate the average values and standard deviations of all expert scores, with results shown in TableĀ 14.
| Stock | Legal | ||||
| Models | GPT-4o | Human expert | Models | GPT-4o | Human expert |
| Qwen-2.5-Instruct | 6.281 | 5.137±0.332 | Qwen-2.5-Instruct | 3.428 | 3.145±0.226 |
| Llama-3-Instruct | 6.129 | 5.025±0.248 | Llama-3-Instruct | 3.417 | 3.136±0.228 |
| OpenO1-Llama | 6.212 | 5.342±0.105 | OpenO1-Llama | 6.554 | 5.808±0.109 |
| OpenO1-Qwen | 6.227 | 5.358±0.133 | OpenO1-Qwen | 6.588 | 5.937±0.121 |
| Finance-LLM | 6.023 | 4.709±0.852 | Open-Australian-Legal | 5.152 | 2.382±0.536 |
| Finance-Chat | 5.583 | 4.832±0.648 | DISC-LawLLM | 0 | 1.403±0.310 |
| Finance-Llama3 | 5.965 | 4.813±0.731 | Law-LLM | 3.838 | 2.820±0.733 |
| FinGPT | 3.413 | 2.101±0.994 | Law-Chat | 3.339 | 3.274±0.384 |
| Llama-2-taiwan-btc | 0 | 1.438±0.455 | Lawma | 0 | 1.255±0.138 |
| SEP | 6.182 | 5.095±0.433 | |||
| Domains-finance | 6.359 | 5.834±0.452 | Domains-Legal | 6.677 | 6.296±0.148 |
The results show that although human expert scores are not completely consistent with GPT-4o, their relative magnitudes are similar. That is, the PROOF score (GPT-4o) assessment of the relative quality of different models is basically consistent with human experts. Additionally, both human experts and GPT-4o give Domains the highest ratings.
Appendix K Application Extension
We demonstrate that the method proposed in this paper suits finance and legal domains and can be extended to service recommendations, risk assessment, biological medicine, and other scenarios where explainable model answers are needed. Domains can provide high-quality, interpretable answers for decision support in these fields.